# Continuous Urban Change Detection from Satellite Image Time Series with Temporal Feature Refinement and Multi-Task Integration

### v2

@@ V-C Ablation study
TFR module: We perform an additional ablation experiment investigating the contribution of the TFR module and testing it with different recurrent sequence models, including recurrent neural networks (RNNs) [60] and LSTMs [29]. Furthermore, we run all settings with and without the MTI module to highlight the complementary nature of the modules. Table IV shows that adding a temporal sequence module improves performances across all three tasks. Among the sequence modules, the transformer encoder is most effective for the SpaceNet 7 dataset. This also holds for the segmentation task on the WUSU dataset. On the other hand, all sequence models achieved similar performances on the WUSU change detection tasks.
TABLE IV: Ablation results for the TFR module with different sequence models on the SpaceNet 7 (SN7) and WUSU datasets. The maximum number of edges was used for each experiment in the change loss. The best and second-best performances are highlighted in red and blue, respectively.
Dataset TFR module Seq. model Change detection Segmentation
@@
SN7 - 0.511 / 0.343 / 98.9 0.344 / 0.208 / 99.6 0.576 / 0.404 / 94.2
RNN 0.525 / 0.356 / 98.9 0.377 / 0.232 / 99.6 0.572 / 0.401 / 93.8
LSTM 0.547 / 0.377 / 99.0 0.402 / 0.252 / 99.6 0.588 / 0.417 / 94.4
TE 0.551 / 0.381 / 99.0 0.414 / 0.261 / 99.7 0.596 / 0.424 / 94.3
WUSU - 0.392 / 0.245 / 96.6 0.339 / 0.205 / 97.6 0.649 / 0.480 / 88.4
RNN 0.440 / 0.282 / 96.9 0.390 / 0.243 / 98.2 0.658 / 0.491 / 88.4
LSTM 0.441 / 0.283 / 96.9 0.387 / 0.240 / 98.2 0.658 / 0.490 / 88.5
TE 0.440 / 0.282 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.496 / 88.9
MTI module: We perform a third ablation experiment investigating performance gains from leveraging additional edges in the MTI module. A degenerate Markov network that uses no change information (i.e., only the segmentation information represented as nodes) is added as a baseline. Table V shows that introducing change information in the MTI module results in considerable change detection performance gains compared to the degenerate setting, as well as minor segmentation performance gains. For example, using the adjacent setting improves change detection performance (first-last) by 44.8 % and 61.6 % in terms of F1 score and IoU, respectively, compared to the degenerate setting on the SpaceNet 7 dataset. On the WUSU dataset, the corresponding performance improvements are 57.5 % and 73.6 % for the F1 score and IoU, respectively (cyclic loss scenario). Table V also shows that introducing change information beyond adjacent edges results in further change detection performance gains, albeit to a much lesser extent. For example, compared to only using adjacent change information, dense information improved the F1 score and IoU values by 2.0 % and 3.0 %, respectively. In the context of the WUSU dataset, additional change information (i.e., adjacent vs. cyclic) improved the F1 score and IoU values by 1.6 % and 2.2 %, respectively. Therefore, the ablation study demonstrates that the proposed MTI module effectively integrates the outputs of the segmentation and change detection tasks at inference time.
TABLE V: Ablation results for the MTI module on the SpaceNet 7 (SN7) and WUSU datasets. The edge settings cyclic and dense are equivalent for the WUSU dataset since its time series consists of three images. The best and second-best performances are highlighted in red and blue, respectively.
@@
Time series length: To investigate the effect of SITS length on performance, we tested the proposed network with different settings for $T$. For the change loss and MTI module, the maximum number of edges (i.e., dense edge setting) for a given $T$ was used (see Equation 6). However, it should be noted that for lengths $T=3$ and $T=2$, the edge settings dense are equivalent to cyclic and adjacent, respectively. The results of this experiment (Fig. 10) show that adding intermediate images has a positive impact on bi-temporal change detection performance. Specifically, while the proposed method performed worse for $T=2$ compared to SOTA methods (Table I), our model outperforms these methods, including multi-temporal change detection methods, at longer time series lengths. Therefore, it effectively uses multi-temporal information to improve change detection performance.
Figure 10: The effect of time series length on bi-temporal change detection performance for the SpaceNet 7 dataset. The maximum number of edges was used for each time series length in the change loss and MTI module. Values represent the mean $\pm$ 1 standard deviation of 5 runs.
@@ VI Discussion
Our results highlight the challenging nature of continuously detecting urban changes from high-resolution SITS. Most methods are developed for bi-temporal urban change detection from VHR imagery, and SOTA methods have achieved remarkable results with F1 scores exceeding 0.9 on datasets such as LEVIR-CD [18] and WHU-CD [45]. In contrast, studies focusing on urban change detection from high-resolution imagery (e.g., PlanetScope or Sentinel-2) often fall short of F1 scores of 0.6, even when employing multi-task learning and multi-temporal data [28]. The suite of methods that achieved these impressive performances typically employs transformers to capture long-range contextual information in bi-temporal VHR imagery [18, 19, 20, 17]. Our experiments indicate, however, that improving long-range context is less relevant in high-resolution imagery (Tables I and II). In addition, transformer-based methods rely heavily on pretraining on large datasets, which ScratchFormer successfully mitigates [22]. Furthermore, a recent study suggests that a simple segmentation ConvNet such as U-Net is still a top performer for change detection tasks [61]. We attribute those findings to the capability of modern ConvNets to capture sufficient spatial context in satellite imagery, especially when ground sampling distances are above 1 m. In comparison to the SOTA methods for bi-temporal change detection, our method leverages the self-attention mechanism to model multi-temporal information in SITS. We showed that the proposed TFR module contributes to the network’s representation learning capability, resulting in improved change detection and semantic segmentation performance (Table IV). Therefore, we deem self-attention an effective mechanism for feature refinement in continuous urban change detection networks.
@@
Our work also highlights the need for effective integration approaches in multi-task learning schemes. Specifically, although multi-task learning is commonly applied for change detection [37, 28], existing multi-task studies do not address the integration of the semantic segmentation and change detection outputs. To fill this research gap, we proposed the MTI module that represents segmentation and change predictions using Markov networks to find the optimal built-up area state for each timestamp in a pixel time series. Our results demonstrate that the proposed integration approach improves both tasks, namely the change detection and building segmentation task (Table V). Furthermore, we demonstrate that the proposed approach benefits from integrating dense change information, obtained from predicting changes between all possible combinations of satellite image pairs in a time series, compared to using only adjacent change information (Table V).
@@ VII Conclusion
In this study, we propose a continuous urban change detection method using optical SITS. The proposed method leverages a transformer-based module to temporally refine feature representations extracted from image time series using a ConvNet, while preserving the temporal dimension. Furthermore, a new multi-task integration approach based on pairwise Markov networks is proposed to combine building segmentation and dense urban change information. We demonstrated the effectiveness of our method on the SpaceNet 7 and WUSU datasets, where it outperformed existing bi-temporal and multi-temporal change detection methods and segmentation methods. The ablation study attributes these accuracy gains to the proposed TFR and MTI modules. Therefore, this research demonstrates the potential of high-resolution SITS for the continuous detection of urban changes.

### v3

@@ V-D Ablation study
TSCD first-last 0.270 / 0.157 / 84.9 0.157 / 0.086 / 62.8
adjacent 0.255 / 0.147 / 89.6 0.453 / 0.296 / 95.1 -
cyclic 0.512 / 0.344 / 92.9 0.553 / 0.383 / 96.4 -
dense 0.543 / 0.373 / 93.7 0.573 / 0.402 / 96.6 -
TFR module: We perform an additional ablation experiment investigating the contribution of the TFR module and testing if recurrent sequence models, particularly recurrent neural networks (RNNs) [60] and LSTMs [29], can be considered as alternatives to self-attention. We run all settings with and without the MTI module due to the complementary nature of the modules. Table VI shows that adding the TFR module to our framework achieves large performance gains across all tasks and datasets. Among the sequence models, self-attention outperformed the recurrent sequence models on the segmentation task across all datasets and the continuous urban change detection task on SpaceNet 7 and the TSCD dataset. It only fell short of LSTMs on the WUSU dataset. Finally, for bi-temporal change detection, self-attention outperformed RNNs on all datasets but LSTMs achieved better results before multi-task integration. In general, self-attention is the most effective sequence model, especially in combination with the MTI module. However, LSTMs should be considered for multi-temporal change detection and datasets with few timestamps.
TABLE VI: Ablation results for the TFR module with different sequence models. The maximum number of edges was used for each experiment in the change loss and MTI module. The best and second-best performances on the SpaceNet 7 (SN7), WUSU, and TSCD datasets are highlighted in red and blue, respectively.
Dataset TFR module Sequence model MTI module Change detection Segmentation
@@
SN7 - 0.490 / 0.324 / 98.8 0.314 / 0.186 / 99.6 0.549 / 0.379 / 93.6
- 0.511 / 0.343 / 98.9 0.344 / 0.208 / 99.6 0.576 / 0.404 / 94.2
Recurrent (RNN) 0.523 / 0.354 / 98.8 0.333 / 0.200 / 99.6 0.556 / 0.386 / 93.5
Recurrent (RNN) 0.525 / 0.356 / 98.9 0.377 / 0.232 / 99.6 0.572 / 0.401 / 93.8
Recurrent (LSTM) 0.549 / 0.378 / 98.9 0.351 / 0.213 / 99.7 0.571 / 0.400 / 94.1
Recurrent (LSTM) 0.547 / 0.377 / 99.0 0.402 / 0.252 / 99.6 0.588 / 0.417 / 94.4
Self-attention 0.537 / 0.367 / 98.8 0.397 / 0.248 / 99.7 0.593 / 0.422 / 94.3
Self-attention 0.551 / 0.381 / 99.0 0.414 / 0.261 / 99.7 0.596 / 0.424 / 94.3
WUSU - 0.342 / 0.208 / 96.8 0.364 / 0.223 / 97.9 0.583 / 0.412 / 86.2
- 0.392 / 0.245 / 96.6 0.339 / 0.205 / 97.6 0.649 / 0.480 / 88.4
Recurrent (RNN) 0.423 / 0.268 / 96.7 0.391 / 0.243 / 98.2 0.613 / 0.443 / 87.1
Recurrent (RNN) 0.440 / 0.282 / 96.9 0.390 / 0.243 / 98.2 0.658 / 0.491 / 88.4
Recurrent (LSTM) 0.426 / 0.271 / 96.8 0.394 / 0.245 / 98.3 0.615 / 0.444 / 87.3
Recurrent (LSTM) 0.441 / 0.283 / 96.9 0.387 / 0.240 / 98.2 0.658 / 0.490 / 88.5
Self-attention 0.420 / 0.266 / 96.6 0.391 / 0.243 / 98.3 0.660 / 0.493 / 88.8
Self-attention 0.440 / 0.282 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.496 / 88.9
TSCD - 0.278 / 0.163 / 89.7 0.295 / 0.173 / 93.7 -
Recurrent (RNN) 0.445 / 0.293 / 91.8 0.497 / 0.334 / 95.5 -
Recurrent (LSTM) 0.561 / 0.391 / 93.8 0.565 / 0.394 / 96.6 -
Self-attention 0.543 / 0.373 / 93.7 0.573 / 0.402 / 96.6 -
MTI module: We perform a third ablation experiment investigating performance gains from leveraging additional edges in the MTI module. Here, we apply multi-task integration with different settings to the outputs of our method trained using the change loss with the maximum number of edges and the segmentation loss. The TFR module using self-attention was enabled. A degenerate Markov network that uses no change information (i.e., only the segmentation information represented as nodes) is added as a baseline. Table VII shows that introducing change information in the MTI module results in considerable change detection performance gains compared to the degenerate setting, as well as minor segmentation performance gains. For example, using the adjacent setting improves change detection performance (first-last) by 44.8 % and 61.6 % in terms of F1 score and IoU, respectively, compared to the degenerate setting on the SpaceNet 7 dataset. On the WUSU dataset, the corresponding performance improvements are 57.5 % and 73.6 % for the F1 score and IoU, respectively (cyclic loss scenario). Table VII also shows that introducing change information beyond adjacent edges results in further change detection performance gains, albeit to a much lesser extent. For example, compared to only using adjacent change information, dense information improved the F1 score and IoU values by 2.0 % and 3.0 %, respectively. In the context of the WUSU dataset, additional change information (i.e., adjacent vs. cyclic) improved the F1 score and IoU values by 1.6 % and 2.2 %, respectively. Therefore, the ablation study demonstrates that the proposed MTI module effectively integrates the outputs of the segmentation and change detection tasks at inference time.
TABLE VII: Ablation results for the MTI module on the SpaceNet 7 (SN7) and WUSU datasets. The edge settings cyclic and dense are equivalent for the WUSU dataset since its time series consists of three images. The best and second-best performances are highlighted in red and blue, respectively.
@@
Time series length: To investigate the effect of SITS length on performance, we tested the proposed network with different settings for $T$ on SpaceNet 7. Additionally, we compared self-attention and LSTM for sequence modeling in the TFR module. The maximum number of edges (i.e., dense edge setting) was used across all $T$ settings in the change loss (see Equation 6). The segmentation loss was enabled, whereas the MTI module was disabled. It should be noted that for lengths $T=3$ and $T=2$, the edge settings dense are equivalent to cyclic and adjacent, respectively. Fig. 10 shows the results of this experiment for bi-temporal change detection (Fig. 10(a)), continuous change detection (Fig. 10(b)), and segmentation (Fig. 10(c)). The bi-temporal results show that adding intermediate images improves change detection performance. However, performance saturates at $T=5$ and even decreases for the recurrent sequence model at $T=6$. Segmentation performance generally also increases with time series length until $T=5$. In contrast, the continuous change detection task increases in difficulty with time series length, since the temporal gap between adjacent image pairs in the time series decreases. Consequently, continuous change detection performance tends to decrease at longer time series lengths. Regarding the sequence model comparison, our network achieved better bi-temporal change detection performance using the recurrent model in the TFR module, except for the longest time series length ($T=6$). However, performance differences are below 5.5 % across all time series lengths. On the other hand, self-attention outperformed the recurrent model for the continuous change detection and segmentation tasks. The largest performance gains in both cases were observed for the longest time series lengths (i.e., $T=5$ and $T=6$).
((a)) Bi-temporal change detection
((b)) Continuous change detection
((c)) Segmentation
Figure 10: The effect of time series length and sequence model in the TFR module on network performance (without MTI module) for the SpaceNet 7 dataset. Network performance is evaluated in terms of (a) bi-temporal change detection, (b) continuous change detection, and (c) segmentation. The maximum number of edges was used for each time series length in the change loss. Values represent the mean $\pm$ 1 standard deviation of 5 runs.
@@ VI Discussion
Our results highlight the challenging nature of continuously detecting urban changes from high-resolution SITS. Indeed, most change detection studies focus on bi-temporal urban change detection from VHR imagery. The suite of methods considered SOTA for bi-temporal change detection typically employ the self-attention mechanism popularized by the transformer architecture to capture long-range contextual information in VHR imagery [18, 19, 20, 17]. On popular benchmark datasets such as LEVIR-CD [18] and WHU-CD [42], these SOTA methods have achieved remarkable results with F1 scores exceeding 0.9. Although transformer-based methods generally outperformed their bi-temporal ConvNet-based counterparts in our experiments, they rarely achieved F1 scores higher than 0.4 on the continuous urban change detection task (Tables II, III, and IV). Therefore, our results indicate that bi-temporal change detection methods generally face limitations for continuous urban change detection.
In comparison to the SOTA methods for bi-temporal change detection, our method leverages the self-attention mechanism to model multi-temporal information in SITS. We showed that the proposed TFR module contributes to the network’s representation learning capability, resulting in improved continuous change detection performance (Table VI). We also compared our method against multi-temporal methods using self-attention or recurrent models for temporal modeling. However, these methods collapse the temporal dimension of the SITS, limiting them to the detection between the first and last image. Although multi-temporal methods generally outperformed bi-temporal methods on this task, they fell short of the proposed method (Tables II, III, and IV). Overall, we deem self-attention an effective mechanism for temporal modeling of SITS. Despite that, our ablation results show that recurrent sequence models such as LSTMs could be considered as an alternative, especially at shorter time series lengths (Table VI).
@@
Our work also highlights the need for effective integration approaches in multi-task learning schemes. Specifically, although multi-task learning is commonly applied for change detection [36, 28], existing multi-task studies do not address the integration of the semantic segmentation and change detection outputs. To fill this research gap, we proposed the MTI module that represents segmentation and change predictions using Markov networks to find the optimal built-up area state for each timestamp in a pixel time series. Our results demonstrate that the proposed integration approach improves both tasks, namely the change detection and building segmentation task (Table VII). Furthermore, we demonstrate that the proposed approach benefits from integrating dense change information, obtained from predicting changes between all possible combinations of satellite image pairs in a time series, compared to using only adjacent change information (Table VII).
Although we demonstrated the effectiveness of our method across three datasets, performances vary significantly across them. Overall, the lowest change detection performances were obtained for the WUSU dataset. The visualizations of the Gaofen-2 SITS highlight several challenging aspects for change detection in this dataset (Figures 8 and 9). In particular, the images were acquired under different atmospheric conditions, and they have large off-nadir angles and contain shadows. In comparison, these artifacts are not apparent in the SpaceNet 7 images (Figures 6 and 7), reducing the complexity of the change detection task. However, building segmentation performance on the SpaceNet 7 dataset is lower than on the WUSU dataset. We attribute this to the fact that individual buildings are more difficult to distinguish in PlanetScope imagery due to its lower spatial resolution (see Table I). The TSCD dataset stands out for its strong continuous urban change detection performances, while performances on the bi-temporal task are lower, especially for bi-temporal change detection methods. Here, it should be considered that the dataset only provides labels between adjacent images. Therefore, we had to derive change labels for non-adjacent images (see Section IV-A), which could affect the reference data quality.
@@ VII Conclusion
This study introduces a continuous urban change detection framework for optical SITS. The proposed method incorporates a transformer-based module to temporally refine feature representations extracted from image time series using a shared ConvNet. Unlike existing temporal modules in multi-temporal change detection methods, our module preserves the temporal dimension, enabling the detection of continuous changes. Additionally, we propose a novel multi-task integration (MTI) approach based on pairwise Markov networks, effectively combining building segmentation and dense urban change information. We evaluated our method on three SITS change detection datasets: SpaceNet 7, the WUSU dataset, and the TSCD dataset. The proposed method outperformed existing bi-temporal and multi-temporal change detection methods and segmentation methods. In particular, our findings show the limitations of bi-temporal methods in continuous change detection, as they cannot fully exploit multi-temporal information in SITS. While multi-temporal change detection methods overcome this limitation, they remain constrained to detecting changes between the first and last images in SITS. Our ablation study further demonstrates the effectiveness of the TFR module in modeling multi-temporal information and the benefits of incorporating dense change information during training. Moreover, it confirms that the MTI module successfully integrates segmentation and change outputs, leading to improved accuracy across both tasks. In summary, this research underscores the potential of high-resolution SITS for continuous urban change detection. Future work will explore weakly supervised and self-supervised change detection methods for SITS to reduce dependence on annotations.